# Reranking Social Media Feeds: A Practical Guide for Field Experiments

### v1

@@ 5.4. Notifications
Throughout the study, researchers must establish channels of communication with the participants. Participants may need guidance, reminders, or feedback to ensure they engage in the study successfully. For example, in a longitudinal study where it is crucial to visit the social media platform regularly, the study design can employ emails to remind the participants that they are enrolled in the study or to acknowledge that their contribution has been recorded. Similarly, when the study is complete, users may receive a post-survey, instructions to remove the extension, or debriefing notes. This step must be handled carefully to comply with Institutional Review Board (IRB) standards for data protection.
Alternatively, user messages can be delivered directly on the platform by editing the DOM of the social media page. The browser extension can add these messages as in-page banners or modal windows, limiting the need to collect user contact information. In the case of post-surveys showing a pop-up or in-page banner when participants visit the social media platform at the end of the study, researchers can prompt immediate and contextual responses, potentially improving the quality of the data collected.
@@ 5.5. Privacy
Web extensions offer a great opportunity to test interventions that social media platforms would never try because of company priorities or because they do not align with the business objective. Nevertheless, despite offering more freedom to researchers, they must commit to ethical standards in designing the experiments and handling the data. Independent oversight, such as Institutional Review Boards (IRBs), should review and approve the study protocol to ensure it meets ethical standards. Participants must be aware of the potential risks and be informed about the data the web extension collects. Designing a registration flow that forces the participants to accept a consent form before collecting any data is essential. It is also crucial to inform participants that their participation is voluntary and that they can withdraw without penalty. If the experiment collects users’ data, it is good practice to implement measures to ensure confidentiality and store the data in a secure database. Depending on where the study is conducted, researchers must comply with local regulations, such as the General Data Protection Regulation (GDPR) in the European Union. Unless specified in the consent form, at the end of the study, we recommend that the participants be guided in uninstalling the extension, or the extension should have a mechanism to turn it off. This step ensures that the researchers do not collect data without the participant’s consent and that the unused extension does not create potential conflicts with other studies the participant may want to join.
6. Conclusion
@@
Social media platforms have recently become less collaborative in supporting independent academic research. This guide offers some practical recommendations on designing field experiments that, using browser extensions, do not require direct involvement of these platforms. Our recommendations are focused on social media feeds, but many aspects generalize beyond this setup. With this guide, we hope to foster a community of independent researchers who aim to investigate social media’s impact on society and contribute to designing healthier online spaces.
@@ Acknowledgements
Our research was supported by the Hoffman-Yee Research Grants at the Stanford Institute for Human-Centered Artificial Intelligence (HAI), the Swiss National Science Foundation, and the Office of Naval Research.

### v2

@@ 5.4. Notifications
Throughout the study, researchers must establish communication channels with participants. Participants may need guidance, reminders, or feedback to ensure they engage in the study successfully. For example, in a longitudinal study where it is crucial to visit the social media platform regularly, the study design can use emails to remind participants that they are enrolled in the study or to acknowledge that their contributions have been recorded. Similarly, upon completion of the study, users may receive a post-survey, instructions to remove the extension, or debriefing notes. This step must be handled carefully to comply with Institutional Review Board (IRB) standards for data protection.
Alternatively, user messages can be delivered directly on the platform by editing the social media page’s DOM. The browser extension can add these messages as in-page banners or modal windows, limiting the need to collect user contact information. In the case of post-surveys showing a pop-up or in-page banner when participants visit the social media platform at the end of the study, researchers can prompt immediate and contextual responses, potentially improving the quality of the data collected.
@@ 5.5. Privacy
Browser extensions offer a great opportunity to test interventions that social media platforms would never try because of company priorities or because they do not align with the business objective. Nevertheless, even with greater freedom, researchers must still commit to ethical standards in designing experiments and handling data. Independent oversight, such as Institutional Review Boards (IRBs), should review and approve the study protocol to ensure it meets ethical standards. Participants must be aware of the potential risks and be informed about the data the browser extension collects. Designing a registration flow that requires participants to sign a consent form before collecting any data is essential. It is also crucial to inform participants that their participation is voluntary and that they can withdraw without penalty. If the experiment collects users’ data, it is essential to implement measures to ensure confidentiality and store the data in a secure database. Depending on where the study is conducted, researchers must comply with local regulations, such as the General Data Protection Regulation (GDPR) in the European Union.
5.6. End of the experiment
Unless specified in the consent form, at the end of the experiment, it is critical to implement an offboarding mechanism to ensure the browser extension stops collecting data and applying interventions. We recommend that participants be guided to uninstall the extension or that the extension include a mechanism to turn itself off. To inform participants that the extension may still be installed, we recommend adding a clear visual signal, such as an overlay banner, with explicit instructions for uninstallation. This step ensures that the researchers do not collect data without the participant’s consent and that the unused extension does not create potential conflicts with other studies the participant may want to join.
6. Discussion
Ethical considerations. Given the central role of social media platforms in contemporary society, it is essential to maintain an independent and critical understanding of their societal impact. Previous research has argued that such inquiry may be ethically necessary to protect democratic processes (Fiske, 2022).
Simultaneously, the approach outlined in this paper enables interventions that can significantly alter participants’ information exposure. By inserting new posts into the feed or modifying existing content, these systems may influence users’ emotions, attitudes, and perceptions. Consequently, researchers must rigorously evaluate the ethical implications of their interventions.
In addition to the privacy concerns previously addressed, researchers should ensure that the content presented to participants reflects a meaningful risk-benefit trade-off. Interventions must be designed to minimize potential harm while providing clear societal or scientific value. As these studies involve human participants and are conducted within real-world information environments, all interventions require review and approval by the appropriate Institutional Review Board (IRB) and must be implemented only with informed participant consent.
Finally, after the experiment ends, we encourage minimizing harm by following established ethical standards for participant debriefing, clearly explaining the study’s purpose, and disclosing any manipulation or deception used in the design.
Limitations. The proposed framework enables researchers to conduct field experiments directly on live social media platforms, thereby enhancing ecological validity. However, this approach has some limitations.
First, the implementation depends on the stability of platform APIs, which may change over time without warnings and disrupt the intervention. Although in our experience, the frontend APIs remain relatively stable for long periods, it is advisable for researchers to implement diagnostic and monitoring tools to ensure that the extension continues to function as intended, especially during extended deployments.
Second, researchers should consider potential feedback loops between the intervention and the platform’s recommendation algorithm. Down-ranking specific types of content and reducing related engagement signals may prompt the algorithm to further limit recommendations of that content. In some instances, this effect may support the intervention’s objectives by reinforcing the down-ranking of undesirable content. However, such algorithmic adaptation can complicate causal interpretation, so researchers should carefully assess how these dynamics may influence their conclusions.
Finally, the current browser extension implementation has been evaluated exclusively on desktop devices. This limitation may introduce biases in both the study population and observed behaviors. Participants may still access the platform on mobile devices, where the intervention is not applied, potentially diluting treatment exposure. This issue can be partially addressed by reminding participants to avoid using other devices during the study period. Additionally, researchers may ask participants whether they accessed the platform through unsupported devices, or estimate this by comparing the recorded user traces to their overall public activity (e.g., comparing posts liked through the extension with all posts the participant liked), and incorporate such information into their analysis. It is important to clearly communicate to participants that accessing the platform through other means will not affect their participation or compensation.
Beyond desktop. Implementing our framework on mobile devices presents unique challenges. Unlike desktop browsers, modifying official mobile applications may violate the platform’s terms of service. Additionally, security features such as certificate pinning prevent interception or modification of network traffic, even on rooted or jailbroken devices.
Stricter app store policies also prevent the distribution of unauthorized versions of major social media applications. Consequently, developing a generic and reproducible mobile solution remains difficult. Researchers interested in extending this framework to mobile environments could consider developing custom applications that use embedded web views to load the mobile version of the platform. Because such custom apps offer greater control over the web view, they can inject largely the same JavaScript logic used in the desktop browser extension. Compared to browser-based extensions, an additional challenge lies in distribution. On Android, custom applications can typically be distributed as APK files outside the official app store. On iOS, however, distribution policies are significantly more restrictive. A possible approach is to use a beta distribution through TestFlight, though Apple’s review process may still reject such applications. Some mobile browsers, such as Microsoft Edge for mobile [^2] https://play.google.com/store/apps/details?id=com.microsoft.emmx.canary, offer limited browser extension support. However, at the time of writing, this feature is only available in beta versions.
@@ 6. Discussion
Conclusion. Social media platforms have recently become less collaborative in supporting independent academic research (Allen and Tucker, 2025). This guide offers practical recommendations for designing field experiments that use browser extensions and do not require direct involvement from these platforms. Our recommendations are focused on social media feeds, but many aspects generalize beyond this setup. With this guide, we hope to foster a community of independent researchers who aim to investigate social media’s impact on society and to contribute to the design of healthier online spaces.
@@ Acknowledgements
This work was supported in part by the National Science Foundation under awards IIS-2403433, IIS-2403434, IIS-2403435, and BCS-2214203; the Swiss National Science Foundation under award P500PT-206953; and a Hoffman-Yee grant from the Stanford Institute for Human-Centered Artificial Intelligence.